# Sense4FL: Vehicular Crowdsensing Enhanced Federated Learning for Autonomous Driving

### v1

@@ III-B The Federated Learning Procedure
denotes the local loss function for street block $b$, $l_{b}$ is the weighting factor with $\sum_{b=1}^{B}l_{b}=1$, and $f(\mathbf{w},\mathbf{x}_{b}^{i})$ denotes the loss function for samples of class $i$ in street block $b$. For object classification/detection tasks in AD, the local loss function for street block $b$ can be cross-entropy loss[18], logistic regression, or a combination of them [43]. Moreover, the weighting factor can be determined according to the vehicle traffic density (i.e., how many vehicles will encounter the situation) or accident probability. Besides, when there are stringent road safety requirements, like in a school zone, a higher weighting factor can be assigned. Without loss of generality, we assume that samples have the same weighting factor if they are in the same street block. Our framework, however, can be easily extended to various weighting settings.
TABLE I: Summary of Important Notations
@@
To obtain the desired global model $\mathbf{w}$, the FL server selects vehicles to participate in each round of the FL training. Let $\mathcal{V}^{(k)}$ denote the available set of vehicles in the $k$-th round, with the cardinality being $V^{(k)}$. Also, we define the vehicle selection decision variable as $a_{v}^{(k)}\in\{0,1\}$, where $a_{v}^{(k)}=1$ indicates vehicle $v$ is selected in round $k$ and $a_{v}^{(k)}=0$ otherwise. After selection, the FL server broadcasts the current global model $\mathbf{w}^{(k)}$ to the selected vehicles. If vehicle $v$ collects the training data in its $m$-th trajectory, the local loss function is given by
@@
where $\mathbf{w}_{v,m}^{(k)}$ is the local model of vehicle $v$ on its $m$-th trajectory in round $k$ and $\mathbf{x}_{v,m}^{i}$ is the $i$-class dataset collected by vehicle $v$ in the $m$-th trajetory.
Each selected vehicle updates its local model by performing $T$ steps of the local stochastic gradient descent (SGD) update  [44][45]. Assuming vehicle $v$ chooses its $m$-th trajectory, the local update at step $t$ can be computed via
@@
where $\eta$ is the learning rate. At the end of the $k$-th round, each selected vehicle obtains the updated local model $\mathbf{w}_{v,m}^{(k),T}$ and sends it to the nearby BS.
The FL server aggregates local models successfully uploaded by vehicles within the time constraint. In global round $k$, the resulting model uploaded by vehicle $v$ is hence given by
@@
and $q_{v,m}^{(k)}$ is the probability that vehicle $v$ follows the $m$-th trajectory in round $k$ and
@@
where $q_{v,m}^{(k),\mathrm{rcv}}$ denotes the probability of successful reception of the local trained model from vehicle $v$ in the $m$-th trajectory, and we will show how to derive it in Section V. Consequently, the global FL model aggregation can be written as
@@
For readers’ convenience, the important notations in this paper are summarized in Table I.
@@
Figure 2: The workflow of Sense4FL.
III-C The Overview of Sense4FL
@@ IV Convergence Analysis of Sense4FL
We first introduce four widely used assumptions on loss function $\mathcal{F}(\cdot)$[18],  [46][47]:
@@
$\Omega^{(k)}\triangleq$ $\delta\sum_{v=1}^{V^{(k)}}\frac{a_{v}^{(k)}\rho_{v}^{(k)}}{\sum_{v=1}^{V^{(k)}}a_{v}^{(k)}\rho_{v}^{(k)}}\underbrace{\sum_{m=1}^{M_{v}^{(k)}}\overline{\xi_{v,m}^{(k)}}\sum_{i=1}^{C}\|p_{v,m}^{i,(k)}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|}_{\text{local divergence}}$ (13)
@@
$=$ $\delta\sum_{v=1}^{V^{(k)}}\frac{a_{v}^{(k)}\rho_{v}^{(k)}}{\sum_{v=1}^{V^{(k)}}a_{v}^{(k)}\rho_{v}^{(k)}}{D_{v,\text{local}}^{(k)}}+{D_{\text{global}}^{(k)}},$
with ${D_{v,\text{local}}^{(k)}}\triangleq\sum_{m=1}^{M_{v}^{(k)}}\overline{\xi_{v,m}^{(k)}}\sum_{i=1}^{C}\|p_{v,m}^{i,(k)}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|$, ${D_{\text{global}}^{(k)}}\triangleq\sum_{i=1}^{C}\|\sum_{v=1}^{V^{(k)}}\frac{a_{v}^{(k)}\rho_{v}^{(k)}}{\sum_{v=1}^{V^{(k)}}a_{v}^{(k)}\rho_{v}^{(k)}}\sum_{m=1}^{M_{v}^{(k)}}\overline{\xi_{v,m}^{(k)}}p_{v,m}^{i,(k)}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|$, $\delta\triangleq\sum_{j=1}^{T-1}(1+\eta\lambda_{\max})^{j}$ with $\lambda_{\max}\triangleq\max_{i}\lambda_{i}$, and $\overline{\xi_{v,m}^{(k)}}\triangleq\frac{q_{v,m}^{(k)}q_{v,m}^{(k),\text{rcv}}}{\sum_{m=1}^{M_{v}^{(k)}}{q_{v,m}^{(k)}q_{v,m}^{(k),\text{rcv}}}}$.
@@
Observation 1. ${D_{v,\text{local}}^{(k)}}$ denotes the local divergence (i.e., the weighted EMD), which measures the divergence between the data distribution of vehicle $v$ and the data distribution of the entire region, and ${D_{\text{global}}^{(k)}}$ denotes the global weighted EMD, which measures the divergence between the data distribution of all selected vehicles and the data distribution of the region. Since $\phi,U$, and $\epsilon$ are independent of our decision variables, the upper bound of training loss is proportional to $\Omega^{(k)}$, which depends on the local divergence and global divergence, i.e., $D_{v,\text{local}}^{(k)}$ and $D_{\text{global}}^{(k)}$, of the selected vehicles. This indicates that minimizing the upper bound of training loss is equivalent to minimizing $\Omega^{(k)}$.
Observation 2. To reduce $D_{v,\text{local}}^{(k)}$, the data distribution of each selected vehicle should be close to the overall data distribution in the region. To reduce $D_{\text{global}}^{(k)}$, the combined data distribution of all selected vehicles should be close to the global regional data distribution. Intuitively, a small local divergence ensures that each local model will not diverge significantly, whereas a small global divergence ensures that all the selected vehicles can collectively train a global model that better represents the region of interest.
Observation 3. $\delta$ is a parameter related to the number of local steps $T$. If a vehicle conducts one SGD step, i.e., $T=1$, then $\delta$ equals 00, yielding $\Omega^{(k)}=D_{\text{global}}^{(k)}$. In this scenario, the process is equivalent to the centralized training. For this reason, only the global divergence $D_{\text{global}}^{(k)}$ matters, i.e., we only need to pay attention to the combined data distribution of all selected vehicles.
@@ V-A Model Reception Probability Analysis
To minimize the upper bound of training loss in (12), we first need to calculate the successful model reception probability $q_{v,m}^{(k),\mathrm{rcv}}$, which is related to latency, i.e., whether a vehicle has sufficient time to upload the model or not. When vehicle $v$ chooses its $m$-th trajectory, it stops data collection after traversing the first ${g}_{v,m}^{(k)}$ street blocks. The selected vehicles need to train the local model and then upload it to the FL server before the deadline for one training round [^4] We assume the delay for model aggregation is negligible, as in [49]..
1) Local model computing: Let $c_{v}$ denote the required number of processing cycles for computing one sample and $f_{v}$ denote the computing capability (in FLOPS) of vehicle $v$. The computing time for vehicle $v$ can be expressed as
@@
2) Local model uploading: We adopt the Orthogonal Frequency Division Multiple Access (OFDMA) scheme for model uploading. Due to the movement of vehicles and their uncertain routes, predicting their channel state information (CSI) accurately is very challenging, if not impossible. As such, we conservatively estimate the uploading time based on the average uploading data rate at the boundary of the cell coverage. The local model uploading time of vehicle $v$ for its $m$-th trajectory in round $k$ is
$t_{v,m}^{(k),\mathrm{up}}=\frac{\omega}{R_{v,m}^{(k),\mathrm{min}}},$ (15)
where $\omega$ is the local model size and $R_{v,m}^{(k),\mathrm{min}}$ is the minimum expected uplink data rate within cellular coverage. Since it is hard to predict the uplink data rate of a vehicle and its associated BS in the future, we assume full cellular coverage and consider the minimum expected data rate within the coverage as a conservative measurement for model uploading time.
The local model should be uploaded before the deadline. Hence, the time allocated for computing and uploading the local model should satisfy
$t_{v}^{(k),\mathrm{comp}}+t_{v,m}^{(k),\mathrm{up}}+t^{\mathrm{trans}}\leq T^{\mathrm{task}},$ (16)
where $t^{\mathrm{trans}}$ represents the time required for the BS to transmit a model to the FL server via a wired link, which is assumed to be a constant. $T^{\mathrm{task}}$ denotes the time budget for one training round. With this time constraint, we can calculate the successful reception probability $q_{v,m}^{(k),\mathrm{rcv}}$ in (9) by
$q_{v,m}^{(k),\mathrm{rcv}}=P(t_{v}^{(k),\mathrm{comp}}+t_{v,m}^{(k),\mathrm{up}}+t^{\mathrm{trans}}\leq T^{\mathrm{task}}).$ (17)
@@ V-B Problem Formulation
In our Sense4FL system, the FL server selects a subset of vehicles and determines when each vehicle should start training after collecting data from a number of street blocks. We define the vector of the vehicle selection index as $\mathbf{a}^{(k)}\triangleq[a_{1}^{(k)},a_{2}^{(k)},...,a_{V^{(k)}}^{(k)}]^{T}$ and the vector of data collection decision variables as $\mathbf{g}_{v}^{(k)}\triangleq[g_{v,1}^{(k)},g_{v,2}^{(k)},\ldots,g_{v,M_{v}^{(k)}}^{(k)}]^{T},\forall v$, where $a_{v}^{(k)}=1$ indicates that vehicle $v$ is selected at round $k$ and $g_{v,m}^{(k)}$ implies that vehicle $v$ stop data collection at the $g_{v,m}^{(k)}$-th street block (or equivalently, collect data from the first $g_{v,m}^{(k)}$ blocks) in its $m$-th trajectory. To minimize the training loss in (12), the optimization problem is formulated as
@@
where $S$ denotes the number of selected vehicles and $c_{v,m}^{(k)}$ represents the number of street blocks from which vehicle $v$ has already collected data at the time of selection. Constraint (18b) restricts the number of selected vehicles due to limited resources, e.g., bandwidth and budget [^5] Since it has been empirically observed that FL performance increases with the number of clients, we use equality to enforce the exact number of vehicle selection.. Constraint (18c) ensures that vehicles stop data collection after the time of vehicle selection and no later than traversing all the street blocks in their trajectories.
@@
$\min_{\mathbf{a},\mathbf{g}_{v}}~{}$ $\delta\sum_{v=1}^{V}\frac{a_{v}\rho_{v}}{\sum_{v=1}^{V}a_{v}\rho_{v}}{D_{v,\text{local}}}+{D_{\text{global}}}$ (19a)
@@
We note that the objective function is the weighted local divergence between the data distribution of each selected vehicle and the data distribution of the region and the global divergence between the combined data distribution of all selected vehicles and the region data distribution, i.e., ${D_{v,\text{local}}}$ and ${D_{\text{global}}}$, which is particularly challenging to solve.
@@ VI-A NP-Hardness of the Problem
The optimization problem we address follows the structure of a non-linear multiple-choice knapsack problem [50, 51], which is widely recognized to be NP-hard and extremely challenging to solve. Compared to its linear counterpart, the complexity of our problem is further exacerbated by the objective function that involves the global divergence between data distribution of selected vehicles and the global data distribution. Since this global divergence results from the combined effect of selected vehicles’ datasets, it is infeasible to decompose the problem into independent subproblems, rendering classic algorithms, such as dynamic programming [52], inapplicable. Besides, the NP-hardness implies that no algorithm can be found to solve the problem in a polynomial time.

### v2

@@ III-B The Federated Learning Procedure
denotes the local loss function for street block $b$, $l_{b}$ is the weighting factor with $\sum_{b=1}^{B}l_{b}=1$, and $f(\mathbf{w},\mathbf{x}_{b}^{i})$ denotes the loss function for samples of class $i$ in street block $b$. For object classification/detection tasks in AD, the local loss function for street block $b$ can be cross-entropy loss, logistic regression, or a combination of them [40]. Moreover, the weighting factor can be determined according to the vehicle traffic density (i.e., how many vehicles will encounter the situation) or accident probability. Besides, when there are stringent road safety requirements, like in a school zone, a higher weighting factor can be assigned [^6] Without loss of generality, we assume that samples have the same weighting factor if they are in the same street block. Our framework, however, can be easily extended to various weighting settings..
TABLE II: Summary of important notations.
@@
III-B1 Local model updating
To obtain the desired global model $\mathbf{w}$, the FL server selects vehicles to participate in each round. Let $\mathcal{V}^{(k)}$ denote the set of available vehicles in the $k$-th round, with the cardinality being $V^{(k)}$. Also, we define the vehicle selection decision variable as $a_{v}^{(k)}\in\{0,1\}$, where $a_{v}^{(k)}=1$ indicates vehicle $v$ is selected in round $k$ and $a_{v}^{(k)}=0$ otherwise. After selection, the FL server broadcasts the current global model $\mathbf{w}^{(k)}$ to the selected vehicles. If vehicle $v$ collects the training data in the trajectory $h_{v,m}$, the local loss function is given by
@@ III-B1 Local model updating
where $\mathbf{w}_{v,m}^{(k)}$ is the local model in round $k$ and $\mathbf{x}_{v,m}^{i}$ is the $i$-class dataset collected by vehicle $v$ in the trajectory $h_{v,m}$.
Each selected vehicle updates its local model by performing $T$ steps of the local stochastic gradient descent (SGD) update [41]. The local update at step $t$ can be computed via
@@
where $\eta$ is the learning rate.
At the end of the $k$-th round, the resulting model uploaded by vehicle $v$ is hence given by
@@
and $q_{v,m}^{(k)}$ is the probability that vehicle $v$ follows the trajectory $h_{v,m}$ in round $k$ and
@@
where $q_{v,m}^{(k),\mathrm{rcv}}$ denotes the probability of successful reception of the local trained model from vehicle $v$, and we will show how to derive it in Section V.
III-B2 Model aggregation
The FL server aggregates local models successfully uploaded by vehicles within the time constraint. Consequently, the global FL model aggregation can be written as
@@ III-B2 Model aggregation
For readers’ convenience, the important notations in this paper are summarized in Table II.
@@
Figure 3: The workflow of Sense4FL framework.
III-C The Overview of Sense4FL Framework
@@ IV Convergence Analysis of Sense4FL
We first introduce four widely used assumptions on loss function $\mathcal{F}(\cdot)$ [42, 43]:
@@
$\Omega^{(k)}\triangleq$ $\delta\sum_{v=1}^{V^{(k)}}\frac{a_{v}^{(k)}\rho_{v}^{(k)}}{\sum_{v=1}^{V^{(k)}}a_{v}^{(k)}\rho_{v}^{(k)}}\underbrace{\sum_{m=1}^{M_{v}^{(k)}}\overline{\xi_{v,m}^{(k)}}\sum_{i=1}^{C}\|p_{v,m}^{i,(k)}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|}_{\text{client divergence}}$ (13)
@@
$=$ $\delta\sum_{v=1}^{V^{(k)}}\frac{a_{v}^{(k)}\rho_{v}^{(k)}}{\sum_{v=1}^{V^{(k)}}a_{v}^{(k)}\rho_{v}^{(k)}}{D_{v,\text{client}}^{(k)}}+{D_{\text{global}}^{(k)}},$
with ${D_{v,\text{client}}^{(k)}}\triangleq\sum_{m=1}^{M_{v}^{(k)}}\overline{\xi_{v,m}^{(k)}}\sum_{i=1}^{C}\|p_{v,m}^{i,(k)}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|$, ${D_{\text{global}}^{(k)}}\triangleq\sum_{i=1}^{C}\|\sum_{v=1}^{V^{(k)}}\frac{a_{v}^{(k)}\rho_{v}^{(k)}}{\sum_{v=1}^{V^{(k)}}a_{v}^{(k)}\rho_{v}^{(k)}}\sum_{m=1}^{M_{v}^{(k)}}\overline{\xi_{v,m}^{(k)}}p_{v,m}^{i,(k)}-\sum_{b=1}^{B}l_{b}p_{b}^{i}\|$, $\delta\triangleq\sum_{j=1}^{T-1}(1+\eta\lambda_{\max})^{j}$ with $\lambda_{\max}\triangleq\max_{i}\lambda_{i}$, and $\overline{\xi_{v,m}^{(k)}}\triangleq\frac{q_{v,m}^{(k)}q_{v,m}^{(k),\text{rcv}}}{\sum_{m=1}^{M_{v}^{(k)}}{q_{v,m}^{(k)}q_{v,m}^{(k),\text{rcv}}}}$.
@@
Observation 1. $D_{v,\text{client}}^{(k)}$} denotes the client divergence (i.e., the weighted EMD), which measures the divergence between the data distribution of vehicle $v$ and the data distribution of the entire region, and $D_{\text{global}}^{(k)}$} denotes the global weighted EMD, which measures the divergence between the data distribution of all selected vehicles and the data distribution of the region. Since $\phi,U$, and $\epsilon$ are independent of our decision variables, the upper bound of training loss is proportional to $\Omega^{(k)}$, which depends on the client divergence and global divergence, i.e., $D_{v,\text{client}}^{(k)}$ and $D_{\text{global}}^{(k)}$, of the selected vehicles. This indicates that minimizing the upper bound of training loss is equivalent to minimizing $\Omega^{(k)}$.
Intuitively, a large client divergence makes FL hard to converge, while a large global divergence makes the distribution of the training dataset different from that of the test dataset. As a result, minimizing the combined objective with these two metrics leads to smaller training loss, which matches our intuition.
Observation 2. To reduce $D_{v,\text{client}}^{(k)}$, the data distribution of each selected vehicle should be close to the overall data distribution in the region. To reduce $D_{\text{global}}^{(k)}$, the combined data distribution of all selected vehicles should be close to the global regional data distribution. Intuitively, a small client divergence ensures that each local model will not diverge significantly, whereas a small global divergence ensures that all the selected vehicles can collectively train a global model that better represents the region of interest.
Observation 3. $\delta$ is a parameter related to the number of local steps $T$. If a vehicle conducts one SGD step, i.e., $T=1$, then $\delta$ equals 0, yielding $\Omega^{(k)}=D_{\text{global}}^{(k)}$. In this scenario, the process is equivalent to the centralized training. For this reason, only the global divergence $D_{\text{global}}^{(k)}$ matters, i.e., we only need to pay attention to the combined data distribution of all selected vehicles.
@@ V-A Model Reception Probability Analysis
To minimize the upper bound of training loss in (12), we first need to calculate the successful model reception probability $q_{v,m}^{(k),\mathrm{rcv}}$, which is related to latency, i.e., whether a vehicle has sufficient time to upload the model or not. When vehicle $v$ follows trajectory $h_{v,m}$ and stops data collection after traversing the first ${g}_{v,m}^{(k)}$ street blocks, it needs to train the local model and upload it to the FL server before the deadline for one training round [^8] We assume the delay for model aggregation is negligible, as in [45]..
1) Local data collection: When vehicle $v$ stops data collection after traversing the first ${g}_{v,m}^{(k)}$ blocks along trajectory $h_{v,m}^{(k)}$, the local data collection time (DCT) $t_{v,m}^{(k),\mathrm{DCT}}$ in round $k$ can be given by
$t_{v,m}^{(k),\mathrm{DCT}}\triangleq\sum_{b\in h_{v,m}({g}_{v,m}^{(k)})}t_{v,b}^{(k)}-\sum_{b\in h_{v,m}({c}_{v,m}^{(k)})}t_{v,b}^{(k)},$ (14)
where $c_{v,m}^{(k)}$ represents the number of street blocks from which vehicle $v$ has already collected data at the time of selection. DCT indicates how long the vehicle spends collecting data after the time of selection.
2) Local model computing: Let $c_{v}$ denote the required number of processing cycles for computing one sample and $f_{v}$ denote the computing capability (in FLOPS) of vehicle $v$. The computing time for vehicle $v$ can be expressed as
@@
3) Local model uploading: We adopt the Orthogonal Frequency Division Multiple Access (OFDMA) scheme for model uploading. Due to the movement of vehicles and their uncertain routes, predicting their channel state information (CSI) accurately is very challenging, if not impossible. As such, we conservatively estimate the uploading time based on the average uploading data rate at the boundary of the cell coverage. The local model uploading time of vehicle $v$ for trajectory $h_{v,m}^{(k)}$ is
$t_{v,m}^{(k),\mathrm{up}}=\frac{\omega}{R_{v,m}^{(k),\mathrm{min}}}+t^{\mathrm{trans}},$ (16)
where $\omega$ is the local model size, $R_{v,m}^{(k),\mathrm{min}}$ is the minimum expected uplink data rate within cellular coverage, and $t^{\mathrm{trans}}$ represents the time required for the BS to transmit a model to the FL server via a wired link, which is assumed to be a constant. Since it is hard to predict the uplink data rate of a vehicle and its associated BS in the future, we assume full cellular coverage and consider the minimum expected data rate within the coverage as a conservative measurement for model uploading time.
The local data collection, local model computing, and uploading should be finished before the deadline, which means
$t_{v,m}^{(k),\mathrm{DCT}}+t_{v}^{(k),\mathrm{comp}}+t_{v,m}^{(k),\mathrm{up}}\leq T^{\mathrm{task}},$ (17)
where $T^{\mathrm{task}}$ denotes the time budget for one training round. With this time constraint, we can calculate the successful reception probability $q_{v,m}^{(k),\mathrm{rcv}}$ in (9) by
$q_{v,m}^{(k),\mathrm{rcv}}=P(t_{v,m}^{(k),\mathrm{DCT}}+t_{v}^{(k),\mathrm{comp}}+t_{v,m}^{(k),\mathrm{up}}\leq T^{\mathrm{task}}).$ (18)
@@ V-B Problem Formulation
In our Sense4FL system, the FL server selects a subset of vehicles and determines when each vehicle should start training after collecting data from a number of street blocks. We define the vector of the vehicle selection index as $\mathbf{a}^{(k)}\triangleq[a_{1}^{(k)},a_{2}^{(k)},...,a_{V^{(k)}}^{(k)}]^{T}$ and the vector of data collection decision variables as $\mathbf{g}_{v}^{(k)}\triangleq[g_{v,1}^{(k)},g_{v,2}^{(k)},\ldots,g_{v,M_{v}^{(k)}}^{(k)}]^{T},\forall v$, where $a_{v}^{(k)}=1$ indicates that vehicle $v$ is selected at round $k$ and $g_{v,m}^{(k)}$ implies that vehicle $v$ stop data collection at the $g_{v,m}^{(k)}$-th street block (or equivalently, collect data from the first $g_{v,m}^{(k)}$ blocks) in its trajectory $h_{v,m}$. To minimize the training loss in (12), the optimization problem is formulated as
@@
where $S$ denotes the number of selected vehicles. Constraint (19b) restricts the number of selected vehicles due to limited resources, e.g., bandwidth and budget [^9] Since it has been empirically observed that FL performance increases with the number of clients, we use equality to enforce the exact number of vehicle selection.. Constraint (19c) ensures that vehicles stop data collection after the time of vehicle selection and no later than traversing all the street blocks in their trajectories.
@@
$\min_{\mathbf{a},\mathbf{g}_{v}}$~ $\delta\sum_{v=1}^{V}\frac{a_{v}\rho_{v}}{\sum_{v=1}^{V}a_{v}\rho_{v}}{D_{v,\text{client}}}+{D_{\text{global}}}$ (20a)
@@
We note that the objective function is the weighted client divergence between the data distribution of each selected vehicle and the data distribution of the region and the global divergence between the combined data distribution of all selected vehicles and the region data distribution, i.e., $D_{v,\text{client}}$} and $D_{\text{global}}$}, which is particularly challenging to solve.
@@ VI-A NP-Hardness of the Problem
The optimization problem we address follows the structure of a non-linear multiple-choice knapsack problem (MCKP) [46], which is widely recognized to be NP-hard and extremely challenging to solve. Compared to its linear counterpart, the complexity of our problem is further exacerbated by the objective function that involves the global divergence between data distribution of selected vehicles and the global data distribution. Since this global divergence results from the combined effect of selected vehicles’ datasets, it is infeasible to decompose the problem into independent subproblems, rendering classic algorithms, such as dynamic programming [47], inapplicable. Besides, the NP-hardness implies that no algorithm can be found to solve the problem in a polynomial time.